When a copper bar of mass m is dropped through a height difference Δh, its potential energy changes by the amount mgΔh, where g is the acceleration of gravity, equal to 9.81 m s-2. Suppose that when the bar hits the ground, all that energy is converted to heat, increasing the temperature of the bar. If the specific heat capacity of the material in the bar is 0.391 J K-1 g-1, calculate the height from which the bar must be dropped to increase the temperature of the bar by 2.68 °C.

__________ m.
1
Expert's answer
2019-03-28T04:12:55-0400

"\Delta h=\dfrac{Cs* \Delta T}{g}=\dfrac{0.391J^{-1}g^{-1}*2.68*10^{3}}{9.81ms^{2}}=106.81m"